OpenStudy (anonymous):

If a person shoots a basketball overhand from a position of 8 feet from the floor, then the path of the basketball through the hoop can be modeled by the parabola y=(-16x^2/0.434v^2)+1.15x+8, where v is the velocity of the ball in ft/sec, y is the height of the hoop and x is the distance away from the hoop. If the basketball hoop is 10 feet high and located 17 feet away, what initial velocity v should the basketball have to go through the hoop? SIDE NOTE: I'm having trouble solving this graphically AND mathematically.

OpenStudy (anonymous):

I don't know I was never that good at basketball.

OpenStudy (anonymous):

this might help if you post this into the physics department but i solved a similar question today using the projectile forumla and looks like you can use the same equation \[y = x \tan(\alpha) - \frac{ g -x^2 }{ 2-v_{0}^2 } *(1 + tg^2 \alpha)\]

OpenStudy (anonymous):

so you have x = 17, and y = 10, put them values in, and solve for v

OpenStudy (anonymous):

Yes, when I tried isolating v, I messed up somewhere. My homework (online) is insisting I solve it graphically, but when I graph the equations y=10 and y=(-4624/0.434x^2)+27.55, my intersect point isn't the same one the online module comes up with. I can't figure out how I might be entering the equation wrong.

OpenStudy (anonymous):

hmm, Sorry :(

OpenStudy (anonymous):

You should get \[ v\approx24.64 \]

OpenStudy (anonymous):

A reasonable speed

OpenStudy (anonymous):

My calculator is giving me a weird intersect point of -.0405858, but the online homework module is saying it's approx. 24.64. I can't figure out where I'm messing up!

OpenStudy (anonymous):

Could either one of you walk me through isolating v?

OpenStudy (anonymous):

first you get \[ 10=27.55\, -\frac{10654.4}{v^2}\\ \frac{10654.4}{v^2}=27.55\, -10= 17.55 \]

OpenStudy (whpalmer4):

Yeah, you just need the parabola to go through (17,10) so \[10 = \frac{-16(17)^2}{0.434v^2} + 1.15(17) + 8\] Subtract 8 from both sides \[2 = \frac{-16(17)^2}{0.434v^2} + 1.15(17)\]Subtract 1.15(17) from both sides\] \[2-1.15(17) = \frac{-16(17)^2}{0.434v^2}\]Multiply both sides by \(0.434v^2\) \[0.434v^2(2-1.15(17)) = -16(17)^2\]Divide both sides by 0.434(2-1.15(17)) \[v^2=\frac{-16(17)^2}{0.434(2-1.15(17))}\]Take square root of both sides \[v = \sqrt{\frac{-16(17)^2}{0.434(2-1.15(17))}}\]

OpenStudy (anonymous):

\[ 10=27.55\, -\frac{10654.4}{v^2}\\ \frac{10654.4}{v^2}=27.55\, -10= 17.55\\ v^2= \frac {10654.4}{17.55}=607.088\\ v= 24.6392 \]

OpenStudy (whpalmer4):

Here's a graph: the horizontal line represents the height of the hoop, and you can see it crosses the curve at x = 17

OpenStudy (anonymous):

How did you enter the equation into the calculator? I'm getting a line whose parabola hits the x axis, not at y=10.

OpenStudy (whpalmer4):

Well, I'll tell you, but it won't do you much good :-) Plot[{10, (-16 x^2)/(0.434 (24.6391)^2) + 1.15 (x) + 8}, {x, 0, 25}, AspectRatio -> Automatic, PlotRange -> {{0, 25}, {0, 15}}]

OpenStudy (whpalmer4):

the equation itself is (-16 x^2)/(0.434 (24.6391)^2) + 1.15 (x) + 8

OpenStudy (anonymous):

Yeah, the homework module was telling me to graph y=10 and then y=(-16x^2)/(0.434x^2)+19.55+8

OpenStudy (anonymous):

subbing x for v

OpenStudy (whpalmer4):

but the parabola does hit the x-axis, of course — it's just at different points. I also graphed y = 10, that was the horizontal blue line.

OpenStudy (whpalmer4):

oh, you don't sub x for v...x and v are independent! you get a different parabola for each different value of v — that's the speed of the ball.

OpenStudy (anonymous):

No, no, I just mean for the purposes of entering the equation into the graph

OpenStudy (anonymous):

there's no v variable..so I used x instead (-4624)/(0.434x^2)-19.55+8

OpenStudy (whpalmer4):

even then. you only get one variable, which is x. you have to enter the value you found for v in place of v. here's a graph with 3 different curves representing our true solution, and v = 23 and v = 26

OpenStudy (whpalmer4):

but x is in the numerator! the number in the denominator is constant for any given parabola of interest. the formula gives us the height of the ball at position x.

OpenStudy (anonymous):

Hrm..okay. But the problem was trying to get me to solve it solely by graphing those two equations. That's why I was having such a hard time. The vertex of the graph I was getting ended at the x axis.

OpenStudy (whpalmer4):

well, you can get it graphically by graphing the equation using x = 17, and picking different versions of v until you find the one that hits the target. but that's a tedious iterative procedure (especially if you don't have a computer to do the graphing for you!)

OpenStudy (whpalmer4):

or you can just solve the parabola, as we did earlier.

OpenStudy (anonymous):

Yeah, just my TI-83.

OpenStudy (anonymous):

Well, thank you! You've both been VERY helpful.